OpenStudy (anonymous):

saidhari bought a toaster oven for $60.this price was 1/4 off the list price what was the list price .

OpenStudy (jdoe0001):

60 * 1/4 = ?

OpenStudy (jdoe0001):

hmmm 1/4 off the list price so 60 is really 3/4 then well

OpenStudy (anonymous):

im not getting what your saying

OpenStudy (jdoe0001):

\(\large { \begin{array}{ccllll} amount& ration \\\hline\\ 60&\frac{4}{4}-\frac{1}{4}\to \frac{3}{4}\\ x&\frac{4}{4}\to 1 \end{array}\implies \cfrac{60}{x}=\cfrac{\frac{3}{4}}{1} }\)

OpenStudy (anonymous):

so would the price be 15

OpenStudy (jdoe0001):

well... it can't be 15... 60 is LESS than the original/list price so it has to be more than 60

OpenStudy (jdoe0001):

\(\bf \cfrac{60}{x}=\cfrac{\frac{3}{4}}{1}\implies \cfrac{60}{x}=\cfrac{3}{4}\implies x=?\)

OpenStudy (anonymous):

what about 240

OpenStudy (jdoe0001):

240 is 60 * 4 that's 4 times 60 so 60 would be 1/4 OF the list price not OFF the list price

OpenStudy (anonymous):

so how would you solve this

OpenStudy (jdoe0001):

cross-multiply the fractions \(\bf \cfrac{60}{x}=\cfrac{\frac{3}{4}}{1}\implies \cfrac{60}{{\color{brown}{ x}}}=\cfrac{3}{{\color{brown}{ 4}}}\implies {\color{brown}{ 4}}\cdot 60={\color{brown}{ x}}\cdot 3\implies x=?\)

OpenStudy (anonymous):

so it will be 720

OpenStudy (jdoe0001):

hmmm cross-multiplly, then divide both sides by 3 \(\bf \cfrac{60}{x}=\cfrac{\frac{3}{4}}{1}\implies \cfrac{60}{{\color{brown}{ x}}}=\cfrac{3}{{\color{brown}{ 4}}}\implies {\color{brown}{ 4}}\cdot 60={\color{brown}{ x}}\cdot 3\implies \cfrac{4\cdot 60}{3}=\cfrac{x\cdot \cancel{ 3 }}{\cancel{ 3 }} \)

OpenStudy (anonymous):

80

OpenStudy (jdoe0001):

yeap